A pharmacist wishes to mix a solution that is 2% minoxidil. She has on hand 50mL of a 1% solution, and she wishes to add some 4% solution to it to obtain the desired 2% solution. How much 4% solution should she add?
Answer by checkley77(12844)

.01*50+.04x=.02(50+x)
.5+.04x=1+.02x
.04x-.02x=1-.5
.02x=.5
x=.5/.02
x=25 ml of 4% is used.
Proof:
.01*50+.04*25=.02(50+25)
.5+1=.02*75
1.5=1.5